OpenStudy (anonymous):

x-y=4 2x+7y=35 Using substitution. Steps would be most helpful :) Thanks.

OpenStudy (radar):

rewrite first as x=4+y now substitute this in second (step 1) 2(4+y) + 7y=35 now solve for y 8 + 2y +7y = 35 9y=35-8 9y=27 y=3

OpenStudy (radar):

solve for x x-3=4 x=7 Was that step by step enough for u aboveit?

OpenStudy (radar):

The value found for y was substituted in the x-y=4

OpenStudy (anonymous):

Thanks for all the help guys :D